Incomplete manufacturing cost data for Ikerd Company for the year are presented as follows for four different situations:

Direct Materials Used Labour Used Manufacturing Overhead Sum of Manufacturing Costs added in period Work in Progress (1/1) Work in Progress (12/31) Cost of Goods Manufactured
1. $127,000 $160,000 $89,000 (a) $33,000 (b) $360,000
2. (C) $180,000 $123,000 $401,000 (d) $40,000 $470,000
3. $80,000 $100,000 (e) $274,000 $60,000 $80,000 (f)
4. $67,000 (g) $75,000 $238,000 $45,000 (h) $270,000


(a) Indicate the missing amount for each letter below. Do not use $ signs or commas when entering your answer:

(a) $Answer

(b) $Answer

(c) $Answer

(d) $Answer

(e) $Answer

(f) $Answer

(g) $Answer

(h) $Answer
Answer #1

Direct material+Direct labor+Overhead=Total manufacturing costs

Total manufacturing costs+Beginning WIP-Ending WIP=Cost of goods manufactured

Direct materials Labor Manufacturing overhead Sum of Manufacturing Costs added in period WIP 1/1 WIP (12/31) Cost of goods manufactured
127000 160,000 89000 (127000+160,000+89000)=$376,000 33000 (376000+33000-360,000)=$49000 360,000
(401000-180,000-123000)=$98000 180,000 123,000 401000 (470,000+40,000-401000)=$109000 40,000 470,000
80,000 100,000 (274000-80,000-100,000)=$94000 274000 60,000 80,000 (274000+60,000-80,000)=$254,000
67000 (238000-67000-75000)=$96000 75000 238000 45000 (238000+45000-270,000)=$13000 270,000